### v1

# fmeffects: An R Package for Forward Marginal Effects

Abstract
Forward marginal effects (FMEs) have recently been introduced as a versatile and effective model-agnostic interpretation method. They provide comprehensible and actionable model explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? We present the R package fmeffects, the first software implementation of FMEs. The relevant theoretical background, package functionality and handling, as well as the software design and options for future extensions are discussed in this paper.
Introduction
Forward marginal effects (FMEs) (Scholbeck et al., 2022) provide simple yet accurate local model-agnostic explanations in terms of forward differences in prediction. They address questions of the form: If we change x by an amount $h$, what is the change in predicted outcome $\widehat{y}$? For instance, given a medical study where a model is trained to predict a patient’s disease risk, FMEs can tell us each patient’s individual change in predicted risk due to losing 5kg in body weight. FMEs thus provide actionable and comprehensible advice for stakeholders, including ones without expertise in machine learning. If the change in predicted risk is substantial enough, doctors may recommend a tailored exercise and nutrition regimen. To avoid interpreting the change in prediction (for instance, the change in predicted disease risk) as a linear effect, Scholbeck et al. (2022) developed a non-linearity measure (NLM) to supplement the FME. Furthermore, in applied statistics, FMEs are typically averaged to an average marginal effect (AME) for an entire data set, which serves as a global (scalar-valued) feature effect measure (Bartus, 2005). However, a single global average does not accurately represent the nuances of a non-linear model or one with complex interactions. As a remedy, Scholbeck et al. (2022) developed the conditional average marginal effect (cAME) estimator for subgroups of observations. They suggest to find subgroups with similar FMEs, for instance, by recursive partitioning (RP). FMEs therefore represent a means to explain models on a local, regional, and global level.
Contributions: We present the R package fmeffects, the first software implementation of FMEs, including the NLM and the cAME. The user interface only requires a pre-trained model and an evaluation data set. The package is designed according to modular principles and thus, various parts of the package are easily maintainable and extensible with additional functionality. This paper introduces the relevant theoretical background, explains the handling of the package with an applied interpretation use case, and discusses the software design and options for future maintenance and extensions.
Related work: Marginal effects (MEs) have a long history in applied statistics and the Stata programming language (StataCorp, 2023). Initially implemented by Bartus (2005), the margins() command is now fully integrated into Stata and provides comprehensive capabilities for various computations and visualizations (Williams, 2012). MEs are typically defined in terms of derivatives of the model w.r.t. a feature. For instance, this variant is the default approach to interpret models in econometrics (Greene, 2012). A less commonly used definition uses forward differences in prediction with varying step sizes (Scholbeck et al., 2022; Mize et al., 2019). Note that - as opposed to forward differences - derivatives are not suitable to explain piecewise constant prediction functions such as tree-based models. In recent years, MEs have gained traction in the R community. The R package margins (Leeper, 2018) was the first port of Stata’s margins() command to R. Other packages related to MEs include ggeffects (Lüdecke, 2018) and marginaleffects (Arel-Bundock, 2023). Our package, fmeffects, mainly differs from such packages in two aspects:
  • Model-agnostic: Other R packages concerned with MEs directly implement support for specific models in R, which typically does not include model classes such as random forests, support vector machines, or neural networks. In contrast, we choose a model-agnostic approach. For one, we provide support for a plethora of models trained via mlr3 (Lang et al., 2019), tidymodels (Kuhn and Wickham, 2020), and caret (Kuhn, 2008); and second, we implement an adapter pattern (Gamma et al., 1994) (see section on Design and options for extensions) that allows extensions with new model classes without substantially interfering with the remaining source code.
  • Implementing new theory surrounding FMEs: The fmeffects package is the first complete software implementation of the theory surrounding novel model-agnostic FMEs as introduced by Scholbeck et al. (2022).
Methodological background
We first provide a brief overview of the relevant theory. For an in-depth introduction, we refer the reader to the theoretical paper by Scholbeck et al. (2022). FMEs can be used for model explanations on the local, regional (also referred to as semi-global), and global level. These differ with respect to the scope of the feature space where the explanation applies. The local level explains models for single data points; the regional level for a subspace of the entire feature space (which, for randomly sampled data, translates to subgroups of observations), and the global level for the entire feature space (which, for randomly sampled data, translates to the entire data set). Increasing the “topographical” scope of the explanation requires increasing amounts of aggregations of local explanations (see for instance, the illustration by Scholbeck et al. (2020) of aggregations of local explanations to global ones for various methods). This can be problematic for non-linear models where local explanations can be highly heterogeneous due to interactions and non-linear effects estimated by the model.
Notation
[…]
Forward marginal effects
The FME can be considered a basic, local unit of interpretation. Given an observation $\bm{x}$, it tells us how the prediction changes if we move $\bm{x}$ to a different location in feature space $\bm{x}^{\ast}$. This change in feature values typically only concerns a subset of features S, where $\bm{x}^{\ast}=(\bm{x}_{S}+\bm{h}_{S},\bm{x_{-S}})$:
$\text{FME}_{\bm{x},\bm{h}_{S}}$ $=\widehat{f}(\bm{x_{S}}+\bm{h}_{S},\bm{x_{-S}})-\widehat{f}(\bm{x})\quad\quad\text{for continuous features $\bm{x_{S}}$}$
For categorical features, this shift is only possible for observations whose observed values differ from the reference location. The observation-specific categorical FME with reference category $c_{j}$ corresponds to:
$\text{FME}_{\bm{x},c_{j}}=\widehat{f}(c_{j},\bm{x_{-j}})-\widehat{f}(\bm{x})\quad\quad\text{for categorical $x_{j}$}$
Non-linearity measure
For continuous features, we can consider $\bm{x}_{S}+\bm{h}_{S}$ a continuous transition of feature values. The associated change in prediction may be misinterpreted as a piecewise linear effect, which is counteracted by the non-linearity measure (NLM). The NLM corresponds to a continuous coefficient of determination $R^{2}$ between prediction function and the linear secant that intersects $\bm{x}$ and $\bm{x}^{\ast}$ (see Fig. 1). We first parameterize this continuous transition through the feature space as a fraction $t\in[0,1]$ of the multivariate step size $\bm{h}_{S}$:
$\gamma(t)=\bm{x}+t\cdot(\bm{h}_{S},0,\dots,0)$
Given the linear secant $g_{\bm{x},\bm{h}_{S}}(\gamma(t))$, the NLM corresponds to:
[…]
Fig. 2 illustrates the setting for multivariate feature changes. The NLM can be approximated via numerical integration, for instance via Simpson’s rule.
Figure 1: Illustration by Scholbeck et al. (2022) of a univariate FME (blue bar) given the prediction function (yellow) and linear secant (green). The NLM indicates how well the secant can explain the prediction function - which is inversely proportional to the integral visualized in red - in relation to how well the most uninformative baseline model - the average prediction - can explain the prediction function.
Figure 2: Illustration of the multivariate NLM by Scholbeck et al. (2022). Left: An exemplary bivariate prediction function and two points to compute an FME. Our observation is located at point (-5, -5), and we create the shortest path through the feature space to reach point (5, 5), which consists of directly proportional changes in both features. Above the path, we see the linear secant and the non-linear prediction function. Right: The multivariate change in feature values can be parameterized as a percentage $t$ of the step size $\bm{h}_{S}$. As a result, the deviation between prediction function and linear secant, as well as the deviation between the prediction function and mean prediction both correspond to a line integral which can be numerically approximated, for instance via Simpson’s rule.
The NLM indicates how well the linear secant is able to explain the prediction function, compared to the baseline model of using the mean prediction. A value of 1 indicates perfect linearity, where the linear secant is identical to the prediction function. For a value of 0, the mean prediction is able to explain the prediction function as equally well as the secant. For negative values, the mean prediction better explains the prediction function than the linear secant (severe non-linearity).
It is, therefore, easiest to interpret FMEs with NLM values close to 1. Although every FME always represents the exact change in prediction, an FME with a low NLM value does not fully describe the behavior of the model in that specific locality. In contrast, an FME with an NLM close to 1 is a sufficient descriptor of the (linear) model behavior. In other words, the NLM serves as an auxiliary diagnostic tool, indicating trust in how well the FME describes the local change in prediction.
Conditional average marginal effect
In order to receive a global model explanation akin to a beta coefficient in linear models, local FMEs can be averaged to a global average marginal effect (AME). However, this does not accurately represent the nuances of a non-linear prediction function. Given a data set $\mathcal{D}$, the conditional average marginal effect (cAME) estimator applies to a subgroup of observations $\mathcal{D}_{[\;]}$:
$\;\text{cAME}_{\mathcal{D}_{[\;]},\bm{h}_{S}}$ $=\widehat{\mathbb{E}_{\bm{X}_{[\;]}}\left[\text{FME}_{\bm{X}_{[\;]},\bm{h}_{S}}\right]}$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{f}\left(\bm{x_{S}}^{(i)}+\bm{h}_{S},\bm{x_{-S}}^{(i)}\right)-\widehat{f}\left(\bm{x^{(i)}}\right)\right)\,\forall i:\bm{x^{(i)}}\in\mathcal{D}_{[\;]}$
Although this estimator can be applied to arbitrary subgroups, an interpretation task demands searching for subgroups with homogeneous local model explanations. Desiderata for such subgroups include within-group effect homogeneity, between-group effect heterogeneity, full segmentation, non-congruence, confidence, and stability. A viable option to partition $\mathcal{D}$ is to run a recursive partitioning algorithm on $\mathcal{D}$ with FMEs as the target which aims at finding subgroups where FMEs are similar. For instance, in fmeffects, both rpart (Therneau and Atkinson, 2019) and ctree() from partykit (Hothorn and Zeileis, 2015) are supported to find subgroups.
User interface and package handling
Local explanations
The fme() function is the central user interface, which - apart from certain parameters - only requires a pre-trained model and a data set (see section Design and options for extensions for details). Further arguments specify the target, feature(s), step size(s), whether to compute NLM values for each FME, and specification of an extrapolation detection method [^3] Model extrapolations are predictions in areas with a low density of training observations. Scholbeck et al. (2022) provide a short introduction to this topic.. The fme() function initiates construction of and computations for a ForwardMarginalEffects object without requiring the user to know R6 (Chang, 2021) syntax.
For this use case, we train a random forest from the ranger package (Wright and Ziegler, 2017) on the bike sharing data set (Fanaee-T, 2013) using mlr3. Note that models trained via tidymodels and caret are also supported. Our goal is to predict and explain the daily bike rental demand in Washington D.C. based on features such as the outside temperature, wind speed, or humidity. We first train the model:
> library(fmeffects)> data(bikes, package = "fmeffects")> library(mlr3verse)> forest = lrn("regr.ranger")> task = as_task_regr(x = bikes, id = "bikes", target = "count")> forest\$train(task)
Then, we simply pass the trained model, evaluation data, and remaining parameters to the fme() function. It returns a ForwardMarginalEffects object which can be analyzed via summary() and visualized via plot() (see Fig. 3). Here, the outside temperature is raised by 5 degrees Celsius ceteris paribus. Each dot represents a single instance. The arrow above the horizontal axis indicates the step size and direction of the feature change in relation to the feature range. The horizontal value of each instance indicates its observed feature value in the provided data set; the vertical axis value is the corresponding FME.
Let us single out the observation with the largest associated FME (brown colored and enlarged in Fig. 3). This observation corresponds to a single day with a recorded temperature of 13.12 degrees Celsius. Increasing the temperature by 5 degrees Celsius on this particular day results in an additional 125 predicted bike rentals. We plot such model explanations for the entire data set and average FMEs to receive a global model explanation. The AME - the global average of FMEs - is 14.5: An increase in temperature by 5 degrees Celsius results in an average change in predicted daily bike rentals by 14.5.
> effects = fme(+   model = forest,+   data = bikes,+   target = "count",+   feature = "temp",+   step.size = 5,+   ep.method = "envelope")> summary(effects)
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5Extrapolation point detection:  envelope, EPs: 81 of 727 obs. (11 %)Average Marginal Effect (AME):  14.4844
> plot(effects)
Figure 3: Plot of univariate FMEs for feature ‘temp’ and step size 5. Each dot represents an observation. The horizontal value represents the observed value of feature ‘temp’. We move each observation’s ‘temp’ value according to the arrow’s direction and length (in this case, 5 units to the right). The respective FME value is indicated by the vertical value of each point. The horizontal bar indicates the AME (the global FME average). An additional smoothing function facilitates interpretations by modeling an approximate pattern of FMEs across the feature range. We single out one observation (brown colored and enlarged) to demonstrate a local interpretation.
Let us take a moment to compare the FME plot with the individual conditional expectation (ICE) (Goldstein et al., 2015) and partial dependence (PD) (Friedman, 2001) plot generated by the R package iml (Molnar et al., 2018). The ICE and PD are one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022). The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case only ‘temp’). The PD is the average of ICEs (in the univariate case, the vertical average) and indicates the global, average prediction when a subset of features is varied for all observations. From Fig. 4, we gather that in contrast to FMEs, the ICE and PD plot is more difficult to interpret. Although we can see a rough trajectory of the feature influence on local and average predictions, it is difficult to pinpoint the exact effects of changing ‘temp’ on the prediction for single observations. Furthermore, ICE curves are more likely to be subject to model extrapolations, a result from predicting in areas where the model was not trained on a sufficient amount of data.
Figure 4: An ICE and PD plot for feature ‘temp’ generated by the R package iml. Each black curve (an ICE) represents predictions for a single instance while only feature ‘temp’ varies. The yellow curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’ on predictions. As opposed to FMEs, this plot is substantially more difficult to interpret.
In many applications, we are interested in interactions of features on the prediction. We diagnosed a univariate effect of ‘temp’ on the predicted amount of bike rentals, but effects may be enhanced or mitigated when other features are varied at the same time. For instance, humidity has a considerable impact on the effects of heat (Russo et al., 2017). We evaluate an increase in temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points (see Fig. 5). For a bivariate change in feature values, the two arrows depict the direction and magnitude of the feature change in the respective variable. Each point is located at the observation’s observed feature values in the provided data set. Its color indicates the FME associated with the bivariate feature change. An increase in the outside temperature by 5 degrees Celsius and a drop in humidity by 10 percentage points is associated with an AME of 19.14. The isolated univariate effect of increasing the outside temperature is boosted by the simultaneous drop in humidity.
> effects.bivariate = fme(model = forest,+               data = bikes,+              target = "count",+              feature = c("temp", "humidity"),+              step.size = c(5, -0.1),+              ep.method = "envelope")> summary(effects.bivariate)
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 82 of 727 obs. (11 %)Average Marginal Effect (AME):  19.1418
> plot(effects.bivariate)
Figure 5: Visualizing bivariate FMEs for an increase in ‘temp’ by 5 degrees Celsius and a simultaneous drop in ‘humidity’ by 10 percentage points. FMEs are highly heterogeneous. We can see mostly positive effects, especially for observations with combinations of lower ‘temp’ and ‘humidity’ values. For various really warm days, the model even predicts a negative effect.
Let us repeat the same procedure as for univariate feature changes and compare the FME plot to an alternative option, the bivariate PD plot (see Fig. 6). As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals, and it does not inform about the distribution of observed feature values, thus not allowing us to evaluate effects of increasing one feature and decreasing another at the same time. Furthermore, the PD plot - like its univariate counterpart - is more likely to be subject to model extrapolations due to predicting for many feature value combinations the model has not seen during training.
Figure 6: A bivariate PD plot (created via the R package iml), one of the few available visualization tools for interactions. We can see the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals. Plugging in larger values for both features ceteris paribus results in more predicted bike rentals on average. As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values, and as a result, we cannot evaluate the effects of increasing one feature and decreasing another at the same time.
Let us now proceed to investigations of non-linearity. Although some sense of non-linearity can be derived from ICE curves (see Fig. 4), it is hard to quantify, and summarizing non-linearity for an ICE with a large variation in the feature of interest would be somewhat meaningless. Furthermore, for bivariate or higher-dimensional changes in feature values, we lose any option for visual diagnoses of non-linearity. In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of their dimensionality:
> effects.bivariate.nlm = fme(model = forest,+                       data = bikes,+                       target = "count",+                       feature = c("temp", "humidity"),+                       step.size = c(5, -0.1),+                       ep.method = "envelope",+                       compute.nlm = TRUE)> plot(effects.bivariate.nlm, with.nlm = TRUE)
Figure 7: Adding NLM computations to the FME plot. Each dot in the left and right plot represents a single observation and its respective FME and NLM value. We can see both linear (bright, yellow colored NLM) and non-linear FMEs (dark, blue colored NLM).
Fig. 7 simply contrasts each observation’s NLM value with its FME value. In this case, we can see both non-linear FMEs (darker, blue colored NLM values) and linear FMEs (brighter, yellow colored NLM values). This is not surprising, considering that the bike sharing data consists of many features, each being capable to influence the effect of changes in ‘temp’ and ‘humidity’ on the prediction.
So far, we have only evaluated changes in continuous features. In many applications, we are often concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process. Note that in spite of the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations. To illustrate this, we switch each non-rainy day’s precipitation status to rainfall. Rainfall has an average isolated effect of lowering daily rentals by 57 bikes (see Fig. 8).
> effects.categ = fme(model = forest,+                data = bikes,+                target = "count",+                feature = "weather",+                step.size = "rain")> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 657 obs. (0 %)Average Marginal Effect (AME):  -57.3322> plot(effects.categ)
Figure 8: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 57 bikes per day.
Regional explanations
In our examples, we can see highly heterogeneous local effects. The more heterogeneous FMEs are, the less information the AME carries. In many practical applications, we are interested in compactly describing the behavior of the predictive model across the feature space, akin to a beta coefficient in a linear model. This is where regional explanations come into play. We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average, but in terms of multiple regional averages (cAMEs).
In fmeffects, this can be achieved by further processing the ForwardMarginalEffects object containing FMEs (and optionally NLM values) using the came() function. This returns a Partitioning object (in this case, an object of the class "PartitioningCTREE", a subclass of the abstract class "Partitioning", see later section on Design and options for extensions).
For the univariate change in temperature by 5 degrees Celsius, we decide to search for exactly 3 subgroups [^4] This value is to be set by the user depending on how many regional explanations are to be found. How many subgroups can be found depends on the data and predictive model. (for a description of this algorithm, see the following section on Design and options for extensions). A summary of the created object informs us about the number of observations, cAME, and standard deviation (SD) of FMEs inside the root node and leaf nodes (the found subgroups). We succeed in finding subgroups with lower SDs while maintaining an appropriate sample size. The root node SD of 22.5 can be successfully split down to 21.5, 8.5, and 17.7 inside the subgroups. By visualizing the tree, we can see how the data was partitioned. Lower temperatures are associated with larger increases in predicted bike rentals when the outside temperature is raised. For outside temperatures lower than 17 degrees Celsius and working days, we can observe the largest cAME of 32 additional bike rentals per day and a cAME of 11 additional daily bike rentals for weekends. For outside temperatures above 17 degrees Celsius, there is only a slight effect of 1.7 additional bikes per day:
> subspaces = came(effects = effects, number.partitions = 3)> summary(subspaces)
PartitioningCtree of an FME objectMethod:  partitions = 3   n      cAME   SD(FME) 646 14.484353 22.499156 * 230 32.265791 21.546565 130 11.172507  8.554322 286  1.689979 17.749573---* root node (non-partitioned)AME (Global): 14.4844
> plot(subspaces)
Figure 9: Finding subgroups of observations with a decision tree. Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME heterogeneity.
Global explanations
When to search for regional explanations thus depends on the heterogeneity of local effects, which can be summarized for the entire model via the ame() function. It uses a default step size of 1 or 0.01 for small feature ranges. For categorical FMEs, it uses every observed category as a reference category. The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points. A large dispersion indicates heterogeneity of FMEs and thus, a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs. A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs, which might be better describable by subgroups and cAMEs, and - if this has been without success - resorting to local model explanations. Recall our example from the previous section on Regional explanations where we split FMEs associated with increasing temperature by 5 degrees Celsius. From the ame() summary, we gather that ‘temp’ has a relatively large SD in relation to its AME, and the interquartile range indicates a large spread of FMEs from -0.47 in the 25% quantile up to 4.74 in the 75% quantile, thus making it a promising candidate to find subgroups with more homogeneous FMEs.
> ame.results = ame(model = forest, data = bikes, target = "count")> summary(ame.results)
Model Summary Using Average Marginal Effects:      feature step size       AME      SD      0.25      0.75   n1      season    spring  -29.3415 30.6189  -38.2318   -6.3963 5482      season    summer    0.5742 21.4782   -8.3499   10.5607 5433      season      fall      9.97 27.4275   -1.9819    30.745 5394      season    winter   13.9861 23.6806    1.0266   23.6969 5515        year         0  -98.6521 67.2816 -156.4247    -19.63 3646        year         1   96.6155 60.1458     22.91  148.1915 3637       month         1    3.9954 13.4096   -1.2981    7.2207 7278     holiday     False   -1.5799 20.5126  -12.5858   14.0711  219     holiday      True  -12.0674 24.2779  -29.8895    6.4375 70610    weekday       Sat  -54.4458 49.1836  -85.8245  -15.5722 62211    weekday       Sun  -78.6803  51.462 -113.7695  -31.2021 62212    weekday       Mon   10.7612 28.3258   -8.7131   30.5779 62313    weekday       Tue   17.9565 24.7443    0.9637   33.5499 62514    weekday       Wed    20.721 22.6949    1.5863   34.3365 62315    weekday       Thu   19.7423 23.9554   -0.2864   34.7076 62416    weekday       Fri    1.6724 35.8582  -23.9194   29.7121 62317 workingday     False -204.2023 90.1427 -256.0286 -141.4579 49618 workingday      True  165.2412 64.8985  123.8688   213.351 23119    weather     clear   25.9084  42.223    3.6162   22.3112 28420    weather     misty    3.3301 33.2101   -9.0864    0.8842 51321    weather      rain  -57.3322 55.5617  -95.7276   -3.8571 65722       temp         1    2.3959  7.2201   -0.4717    4.7369 72723   humidity      0.01   -0.1681  2.4691   -0.2999    0.4329 72724  windspeed         1    0.0198  2.2861   -0.2156    0.2126 727
Design and options for extensions
The fmeffects package is built on a modular design for improved maintainability and future extensions. Fig. 10 provides a visual overview of the core design. Major emphasis is given to the strategy and adapter design patterns (Gamma et al., 1994). Simply put, the strategy pattern decouples source code for algorithm selection at runtime into separate classes. We repeatedly implement this pattern throughout the package by creating abstract classes whose subclasses implement specific functionalities. The adapter design pattern (also referred to as a “wrapper”) creates an interface for communication between two classes.
Figure 10: Design overview of the fmeffects package, including methods that implement the main functionality of each class. Classes may contain more methods than depicted. Blue boxes indicate wrapper functions to instantiate objects of the respective class.
  • "Predictor": An abstract class that implements the adapter pattern to accommodate future implementations of storing a predictive model. "PredictorMLR3", "PredictorParsnip", and "PredictorCaret" are subclasses that store an mlr3, parsnip (Kuhn and Vaughan, 2023) (part of tidymodels), or caret model object. This allows users of fmeffects to use numerous predictive models such as random forests, gradient boosting, support vector machines, or neural networks. The package can be extended with novel model types by implementing a new subclass that stores the model, data, target, and is able to return predictions.
  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset thereof). Internally, a new "ForwardMarginalEffects" object is used to compute FMEs and aggregate them. For convenience, we implement a wrapper function ame() to facilitate object creation and to initiate computations without requiring user input in the form of R6 syntax.
  • "ForwardMarginalEffects": The centerpiece class of the package. It keeps access to a Predictor, stores important information to create FMEs, and after computations are completed, stores results and gives access to visualization methods. For convenience, the wrapper function fme() can be used.
  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotCategorical".
  • "ExtrapolationDetector": Identifies (and excludes) extrapolation points. The current implementation supports the method “envelope”, which excludes points outside of the multivariate envelope of the training data (Scholbeck et al., 2022).
  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, for instance via Simpson’s 3/8 rule. The general definition for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
    $\int_{a}^{b}f(x)\approx\frac{b-a}{8}\left[f(a)+3f\left(\frac{2a+b}{3}\right)+3f\left(\frac{a+2b}{3}\right)+f(b)\right]$ (1)
    We make use of a composite Simpson rule, which divides up the interval $\left[a,b\right]$ into $n$ subintervals of equal size and approximates each subinterval with Eq. (1).
  • "Partitioning": An abstract class, allowing various implementations of finding subgroups for cAMEs. For convenience, the wrapper function came() can be used. The current implementation supports RP via the rpart and partykit (CTREE algorithm) packages (classes "PartitioningRPart" and "PartitioningCTREE"). We believe there are two criteria that should guide this process:
    • FME homogeneity inside each subgroup.
    • The number of subgroups.
    A low number of subgroups is generally preferred. In certain applications, we may want to search for a predefined number of subgroups, akin to the search for a predefined number of clusters in clustering problems. Many RP algorithms do not support the search for a number of subgroups, which is where the "Pruner" class comes into play.
  • "Pruner": To receive a predefined number of subgroups for arbitrary RP algorithms, we follow a two-stage process: Grow a large tree by tweaking tree-specific hyperparameters and then prune it back to receive the desired number of subgroups. A "Partitioning" subclass is implemented such that it can first grow a large tree, for instance, with a low complexity parameter for rpart. Then, "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
  • "PartitioningPlot": Decouples visualizations of separating $\mathcal{D}$ into subgroups from the concrete implementation of the "Partitioning" subclass. Here, we make use of a dependency on partykit for a tree data structure. This allows visualizations of any partitioning with the same methods. The package ggparty (Borkovec and Madin, 2019) creates tree figures that illustrate the partitioning, descriptive statistics for each terminal node, and histograms of FMEs (and optionally NLM values).
Conclusion
This paper introduces the R package fmeffects, the first software implementation of FMEs. We showcase the package functionality with an applied use case and discuss design choices and implications for future extensions. FMEs are a novel and versatile method for model-agnostic interpretations. They give us explanations in the form of: If we change $x$ by an amount $h$, what is the change in predicted outcome $\widehat{y}$? FMEs provide comprehensible and actionable advice for stakeholders, including ones without machine learning expertise. We therefore hope that this package will work towards a more widespread adaption and usage of FMEs in practice.